OpenStudy (anonymous):

two bodies seperated by a distance of s start moving towards each other with speeds of v and 2v respectively the uniform acceleration with which the first body should move so that they meet at the middle?

OpenStudy (anonymous):

the answer is 8v^2/s

OpenStudy (anonymous):

let the time in which the particles in the middle as t then 2v*t=s/2 or t=s/4v now for 2nd particle we have s/2=v*t+1/2*a*t^2 where a is the required acceleration or s=2vt+a*t^2 we know that 2vt=s/2 hence a*t^2=s/2 put the value of t=s/4v here we get a=8V^2/2s

OpenStudy (anonymous):

2v*t=s/2 did nt understand

OpenStudy (anonymous):

the velocity of a particle is given as 2v which is constant throughout hence 2v*t=s/2 this particle moves with constant velocity to reach in the middle to cover up a distance of s/2..(velocity*time=distance)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanzz

OpenStudy (anonymous):

here time is the key which is same for both the particles.. both the particles require the same time to reach in the middle..